OpenStudy (anonymous):

1.A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (saifoo.khan):

5(x+40)=6(x-40) 5x+200=6x-240 x=440 The speed of the plane in still air is 440 mph..